OpenStudy (anonymous):

Help for a medal, anyone? I'm rusty on inequalities and need some help on a few problems to boost my memory. Here's the first one... -9 < 4x + 8 < 10

OpenStudy (amistre64):

what are your thoughts ...

OpenStudy (anonymous):

If I remember correctly, you're supposed to simplify the equation by adding, divining..etc terms on either side, but I'm not sure how to do that with an inequality.

OpenStudy (anonymous):

dividing*

OpenStudy (amistre64):

the math doesnt change ... what you do to one part, you do to all the parts to keep them "equal" right?

OpenStudy (amistre64):

in this case, the equation is in the middle instead of to the left or right ... what would you think we should do first?

OpenStudy (amistre64):

just make a guess at what you think we should do to start with ... what comes to mind? 4x+8 is what we want to manipulate, how do we work about getting 'x' all by itself?

OpenStudy (anonymous):

Would you solve one equality at a time like -9 < 4x + 8 or 4x + 8 < 10 ?

OpenStudy (amistre64):

you can if you want, but i just work it all at once to avoid the extra work of writing it all out twice

OpenStudy (amistre64):

-9 < 4x + 8 < 10 -8 -8 -8 ---------------- -17< 4x < 2

OpenStudy (anonymous):

Ah, I see. Then you would divide all 3 parts by 4 to get x by itself.

OpenStudy (amistre64):

correct, the only thing to watch out for is when we have to mulitply or divide by a negative number ... which they really should include the rule when doing equalities the sign reverses ... but we do not have to worry about that in this particular case do we

OpenStudy (amistre64):

if -a < b, then a > -b if -a > b, then a < -b if -a = b, then a = -b ... the reversing of an equals sign just produces an equals sign, but it helps to make the rule consistent for all 3 equality signs.

OpenStudy (anonymous):

So would -17/4 < x < 1/2 be the final answer or is there another step?

OpenStudy (amistre64):

thats the answer for me, but im not grading it

OpenStudy (amistre64):

17/4 cant be reduced, but it can be written as 4 and 1/4 if need be

OpenStudy (anonymous):

Okay, thank you. :) Could you help with one other? It looks a bit different from the first.

OpenStudy (amistre64):

i spose

OpenStudy (anonymous):

\[\left|\frac{ x+4 }{ 5 } \right| < 6\]

OpenStudy (amistre64):

absolute value, what does that mean to you?

OpenStudy (anonymous):

I know that nothing inside of it can be negative (or at least that I recall)

OpenStudy (amistre64):

first off |5| = 5 so we can go ahead and multiply thru by 5 |x+4| < 30

OpenStudy (amistre64):

it can be negative inside of it, but the bars cancel out the sign

OpenStudy (amistre64):

|-30| = 30 and |30| = 30 when does x+4 = 30? or -30?

OpenStudy (amistre64):

as a side note: |a| < k can be written as: -k < a < k

OpenStudy (amistre64):

|x+4| < 30 can be written as -30 < x+4 < 30

OpenStudy (anonymous):

Then you just subtract 4 from both sides... -34 < x < 26 Also does the |a| < k or -k < a < k rule always apply?

OpenStudy (amistre64):

it is a rule, so it always applies yes. its just my brain that doesnt always remember it in time lol

OpenStudy (amistre64):

\[|\frac{x+4}{5}|<6\] \[-6<\frac{x+4}{5}<6\] \[-306<x+4<30\]

OpenStudy (amistre64):

not -306, but yeah

OpenStudy (anonymous):

Oh, okay. So -34 < x < 26 would be the answer correct?

OpenStudy (amistre64):

looks good to me

OpenStudy (anonymous):

Thank you so much for your help. :)

OpenStudy (amistre64):

youre welcome, and good luck :)